# Efficient & Correct Predictive Equivalence for Decision Trees

### v2

@@ Abstract
The Rashomon set of decision trees (DTs) finds several relevant uses. Recent work showed that DTs computing the same classification function, i.e. predictive equivalent DTs, can represent a significant fraction of the Rashomon set. Such redundancy can be undesirable. For example, feature importance based on the Rashomon set becomes inaccurate due the existence of predictive equivalent DTs, i.e. DTs that make the same prediction for every possible input. In recent work, McTavish et al. proposed solutions for several computational problems related with DTs, including that of deciding predictive equivalent DTs. This approach, which this paper refers to as MBDSR using the authors’ last names initials, consists of applying the well-known method of Quine-McCluskey (QM) for obtaining minimum-size DNF (disjunctive normal form) representations of DTs, which are then used for comparing DTs for predictive equivalence. Furthermore, the minimum-size DNF representation was also applied to computing explanations for the predictions made by DTs, and to finding predictions in the presence of missing data. However, the problem of formula minimization is hard for the second level of the polynomial hierarchy, and there exist different reasons why the QM method may exhibit worst-case exponential running time and space. This paper first demonstrates that there exist decision trees that trigger the worst-case exponential running time and space of the QM method. Second, the paper shows that, depending on the implementation of the QM method used, the MBDSR approach may produce incorrect results for the problem of deciding predictive equivalence. Third, the paper shows that any of the problems to which the minimum-size DNF representation has been applied to can in fact be solved in polynomial time, in the size of the DT. The experiments confirm that, for DTs for which the worst-case of the QM method is triggered, the algorithms proposed in this paper are orders of magnitude faster than the ones proposed by McTavish et al. Finally, the paper relates predictive equivalence with logic-based explanations and measures of importance.
@@ 1 Introduction
In decision tree learning, the Rashomon set (Breiman, 2001b; Xin et al., 2022) captures the set of decision trees (DTs) for a given classification problem which are within some distance from some accepted definition of optimal DT (Xin et al., 2022). Several uses of the Rashomon set have been studied in recent years (Fisher et al., 2019; Semenova and Rudin, 2019; Rudin et al., 2022; Semenova et al., 2022; Xin et al., 2022; Donnelly et al., 2023; Paes et al., 2023; Biecek et al., 2023; Kobylinska et al., 2024; Hsu et al., 2024; Ciaperoni et al., 2024; Li et al., 2024b, a; Donnelly et al., 2025; Nguyen et al., 2025). For the same classification problem, different decision trees capture different classification functions, and it is often of interest to study a sample of the decision trees in the Rashomon set. Recent work (McTavish et al., 2025c, a) studies predictive equivalent DTs in the Rashomon set, i.e. DTs which represent the same classification function. This earlier work observes that the predictive equivalent DTs in the Rashomon set can represent a fairly significant percentage of the DTs, and so their removal enables keeping in the Rashomon set only DTs that capture different classification functions. For example, this yields better estimates in terms of feature importance (Donnelly et al., 2023).
@@
To identify DTs that are predictive equivalent, McTavish et al. (McTavish et al., 2025c, a) propose a solution that revisits work on boolean formula minimization from the 1950s, namely the well-known Quine-McCluskey (QM) boolean formula minimization method (Quine, 1952, 1955; McCluskey, 1956). [^1] There are different algorithms that implement variants of the QM method (McCluskey, 1956; Brayton et al., 1984; Coudert, 1994). The work of McTavish et al. (McTavish et al., 2025c, a) exploits the implementation of QM available in SymPy (Meurer et al., 2017). Starting from some DT $\mathcal{T}$, the QM method is applied to computing a minimum-size boolean formula $\mathcal{T}_{\mathsf{DNF},c}$ for each class $c$, in disjunctive normal form (DNF). This formula is then used not only for deciding predictive equivalence, but also to answer a number of important queries regarding DT prediction in terms of non-availability of some features and their explanation (McTavish et al., 2025c, a). It should be underlined that the formula ${\mathcal{T}}_{\mathsf{DNF},c}$ is not any DNF representation of a DT for predicting class $c$. Since the language of DTs is less expressive that that of DNF formulas (Hyafil and Rivest, 1976), then any DT can be represented by a DNF of size no larger that the DT. For each class $c$, ${\mathcal{T}}_{\mathsf{DNF},c}$ represents a minimum-size DNF formula obtained from the set of prime implicants for predicting that class. Moreover, ${\mathcal{T}}_{\mathsf{DNF}}$ represents the set of DNFs, one for each class $c$. Throughout this paper, we will use the acronym MBDSR to refer to the approach proposed in the work of McTavish et al. (McTavish et al., 2025c, a).
@@
While it is known that the number of prime implicants can grow exponentially with the number of variables for arbitrary boolean formulas (Chandra and Markowsky, 1978), this paper shows that this is also the case with DTs, i.e. there exist DTs for which the BCF is worst-case exponential in the size of the DT. The paper then improves on earlier work (McTavish et al., 2025c, a), by proving that the computation of $\mathsf{BCF}_{c}(\mathcal{T})$ (which is worst-case exponential in size) and $\mathcal{T}_{\mathsf{DNF},c}$ (which requires solving an NP-hard optimization problem on a worst-case exponential size representation) are both unnecessary, and can be replaced with polynomial-time algorithms. Concretely, this paper proves that the properties of: i) completeness (Theorem 3.2 in (McTavish et al., 2025a)); ii) succinctness (Proposition 3.3 in (McTavish et al., 2025a)); and iii) the resolution of predictive equivalence (Theorem 3.4 in (McTavish et al., 2025a)) can all be solved in polynomial time in the size of the given DT. In addition, we prove that all the algorithms proposed in this paper are correct.
We also note that the issues with McTavish et al.’s approach can go beyond the inefficiency of the proposed algorithms (McTavish et al., 2025c, a). As this paper details, the use of the QM method for deciding predictive equivalent as proposed by McTavish et al. hinges on a requirement of determinism of some steps of the QM method, which may or may not be guaranteed by practical implementations. The QM method produces a minimum-size DNF representation, which lacks canonicity (Brown, 1990). To address this issue, McTavish et al. (McTavish et al., 2025c, a) impose a condition on implementations of the QM method, which may not be guaranteed in practice.
@@
This paper is organized as follows. Section˜2 introduces the notation and definitions used throughtout. Section˜3 develops an in-depth analysis of the work of McTavish et al. (McTavish et al., 2025c, a). Section˜3.1 analyzes the MBDSR approach in terms of running time, proving that the MBDSR approach can require exponential time and space, and also demonstrating that the worst-case running time and space is exercised in the case of DTs. Section˜3.2 goes a step further, and shows that MBDSR can produce incorrect results in the case of deciding predictive equivalence. Section˜4 fixes both the issues of correctness and performance of the MBDSR approach, by proposing polynomial-time algorithms for the computational problems studied by McTavish et al. (McTavish et al., 2025c, a). Section˜6 proves a number of additional results, relating predictive equivalence with logic-based explanations (Marques-Silva, 2022; Marques-Silva and Ignatiev, 2022), but also with measures of importance, including the Shapley values (Shapley, 1953; Chalkiadakis et al., 2012) and Banzhaf values (Banzhaf III, 1965; Chalkiadakis et al., 2012). These additional results also allow developing an algorithm for predictive equivalence that asymptotically improves on the algorithm proposed in Section˜4. Section˜5 presents simple experiments demonstrating the massive difference in performance between the MBDSR approach, which can produce incorrect results, and the algorithms proposed in this paper. Section˜7 concludes the paper.
@@
The authors of the MBDSR approach (McTavish et al., 2025c, a) have been informed of the efficiency and soundness issues in their work that the present paper reveals and corrects (Marques-Silva and Ignatiev, 2025).
@@ 2.1 Machine Learning (ML) Models
A classification problem is defined on a set of features ${\mathcal{F}}=\{1,\ldots,m\}$. Each feature $i$ has a domain $\mathbb{D}_{i}$. Domains can be categorical or ordinal. If ordinal, domains can be integer or real-valued. Feature space is defined as the cartesian product of the domains, $\mathbb{F}=\mathbb{D}_{1}\times\cdots\times\mathbb{D}_{m}$. A classifier computes a classification function $\kappa$, that maps feature space to a set of classes $\mathbb{K}=\{c_{1},\ldots,c_{K}\}$.
@@ 2.2 Logic Foundations
Let $\mathbb{OP}$ denote some universe of relational operators, e.g. we could consider $\mathbb{OP}=\mathbb{OP}_{s}\cup\mathbb{OP}_{t}$, with $\mathbb{OP}_{t}=\{\in\}$ and $\mathbb{OP}_{s}=\{=,<,\leq,\geq,>\}$. (Logic negation can be used to obtain additional operators.) Let $\mathbb{L}=\{(x_{i},R_{i},\mathsf{op}_{i})\,|\,(i\in{\mathcal{F}})\land(R_{i}\subseteq\mathbb{D}_{i})\land(\mathsf{op}_{i}\in\mathbb{OP}_{t})\}\cup\{(x_{i},R_{i},\mathsf{op}_{i})\,|\,(i\in{\mathcal{F}})\land(R_{i}\in\mathbb{D}_{i})\land(\mathsf{op}_{i}\in\mathbb{OP}_{s})\}$ denote all the literals that can be obtained from the variables associated with the features in $\mathcal{F}$} and subsets or values from their domains. (Throughout the paper, we will simplify the notation. For example, $(x_{i},v_{i})$ will be used instead of $(x_{i},v_{i},=)$, and both represent the literal $x_{i}=v_{i}$. Moreover, $(x_{i},v_{i},=)$ is equivalent to writing $(x_{i},\{v_{i}\},\in)$.) The set of partial assignments is defined by the powerset of $\mathbb{L}$, i.e. $\mathbb{A}=2^{\mathbb{L}}$. Thus, a partial assignment ${\mathcal{A}}\in\mathbb{A}$ is a subset of $\mathbb{L}$. In the case of complete assignments (or inputs to an ML model), we require that all features are assigned a single value from their domain, $\mathbb{I}=\{{\mathcal{A}}\in\mathbb{A}\,|\,\forall(i\in{\mathcal{F}}).|\{(x_{i},R_{i},\mathsf{op}_{i})\in{\mathcal{A}}\}|=1\land(x_{i},v_{i},=)\in{\mathcal{A}}\}$. With each literal $(x_{i},R_{i},\mathsf{op}_{i})\in\mathbb{L}$ we associate a domain $\mathsf{dom}((x_{i},R_{i},\mathsf{op}_{i}))=\{\mathbf{x}=(x_{1},\ldots,x_{i},\ldots,x_{m})\in\mathbb{F}\,|\,x_{i}\,\mathsf{op}_{i}\,{v_{i}}\}$. Thus, the domain of a literal is the set of points in feature space for which the literal holds true. Similarly, we define the domain of a partial assignment $\mathcal{A}$}, i.e. $\mathsf{dom}({\mathcal{A}})=\cap_{(x_{i},v_{i},\mathsf{op}_{i})}\mathsf{dom}(x_{i},v_{i},\mathsf{op}_{i})$. In addition, a partial assignment is said to be consistent, it if exhibits a non-empty domain, i.e. $\mathsf{Consistent}({\mathcal{A}})=[\mathsf{dom}({\mathcal{A}})\not$$\emptyset$]. By default, we will require partial assignments to be consistent. Finally, we note that (prime) implicants can also be represented as partial assignments (Dillig et al., 2012; Amgoud, 2023). This provides a convenient framework for formalizing explanations in DTs.
@@ 2.3 Classification with Decision Trees
A DT is a connected directed acyclic graph ${\mathcal{T}}=(V,E)$, with a single root node $r\in{V}$. $r=\mathsf{Root}({\mathcal{T}})$ indicates that $r\in{V}$ is the root of DT $\mathcal{T}$}. The tree nodes $V$ are partitioned into non-terminal and terminal nodes. Each non-terminal node $s$ (such that $\mathsf{NonTerminal}(s;{\mathcal{T}})$ holds true) is associated with a feature $i$, i.e. $i=\mathsf{Feature}(s;{\mathcal{T}})$. For a non-terminal node $s$, associated with feature $i$, the domain of $i$ is partitioned among the children of node $s$. Each terminal node $s$ (such that $\mathsf{Terminal}(s;{\mathcal{T}})$ holds true) is associated with a class from $\mathbb{K}$, denoted by $\mathsf{Class}(s;{\mathcal{T}})$. Whereas non-terminal nodes have two or more children, terminal nodes have no children. Moreover, for each node $s$ in the DT, $\mathsf{Parent}(s;{\mathcal{T}})$ represents the parent of $s$ in the tree. For the root node $r=\mathsf{Root}({\mathcal{T}})$, $\mathsf{Parent}(r;{\mathcal{T}})=\mathsf{none}$. As with any classifier, a DT computes a classification function $\kappa_{{\mathcal{T}}}$, mapping feature space to $\mathbb{K}$. For convenience, $\mathsf{NonTerminal}({\mathcal{T}})$ denotes the nodes in $V$ that are non-terminal. Similarly, $\mathsf{Terminal}({\mathcal{T}})$ denotes the nodes in $V$ that are terminal.
@@
Two DTs $\mathcal{T}$} and $\mathcal{T}$$^{\prime}$ are predictive equivalent if it holds that: $\forall(\mathbf{x}\in\mathbb{F}).\left(\kappa_{{\mathcal{T}}}(\mathbf{x})=\kappa_{{\mathcal{T}}^{\prime}}(\mathbf{x})\right)$. Thus, predictive equivalence signifies that the two DTs cannot be distinguished over any point in feature space. For any pair of DTs ${\mathcal{T}}_{1},{\mathcal{T}}_{2}$, the predicate $\mathsf{PredEquiv}({\mathcal{T}}_{1},{\mathcal{T}}_{2})$ takes value $\top$ (i.e. it holds true) if the pair of DTs is predictive equivalent; otherwise it takes value $\bot$. Although the paper focuses mostly on DTs, it is plain that the definition of predictive equivalencd can be used to relate two arbitrary ML models. For example, we could have a DT that is predictive equivalent to random forest (RF) (Breiman, 2001a). In Section˜6 we revisit predictive equivalence in this more general setting.
Although it is well-known that learning an optimal decision tree is NP-hard (Hyafil and Rivest, 1976), recent years witnessed a renewed interest in the learning of optimal decision trees (e.g. (Nijssen and Fromont, 2007; Bessiere et al., 2009; Nijssen and Fromont, 2010; Bertsimas and Dunn, 2017; Verwer and Zhang, 2017; Narodytska et al., 2018; Verwer and Zhang, 2019; Hu et al., 2019; Verhaeghe et al., 2020a; Avellaneda, 2020; Aglin et al., 2020a; Lin et al., 2020; Janota and Morgado, 2020; Hu et al., 2020; Aglin et al., 2020b; Verhaeghe et al., 2020b; Demirovic and Stuckey, 2021; Shati et al., 2021; Demirovic et al., 2022; Rudin et al., 2022; van der Linden et al., 2023), among many others), with the purpose of improving ML model interpretability. The rationale is that DTs are interpretable and so the model can be used for obtaining an explanation, e.g. the path consistent with the given instance (Molnar, 2020). Moreover, and although it is generally accepted that DTs are interpretable (Breiman, 2001b), [^4] The perception of DT interpretability can be traced back more than two decades. A quote from a 2001 paper by L. Breiman (Breiman, 2001b) illustrates this perception: “On interpretability, trees rate an A+”. there is no rigourous definition of interpretability or interpretable ML model and this is unlikely to change (Lipton, 2018). Furthermore, a number of recent works (Izza et al., 2020, 2022; Audemard et al., 2022b; Marques-Silva and Ignatiev, 2023; McTavish et al., 2025a) observe that even optimal decision trees need to be explained, e.g. using a formal definition of explanation (Marques-Silva, 2022; Marques-Silva and Ignatiev, 2022). Concretely, recent work demonstrated that if DTs are explained by reporting the path consistent with some instance, then the explanation can be arbitrarily redundant on the number of features (Izza et al., 2020, 2022). Accordingly, one of the concerns of the MBDSR approach of McTavish et al. (McTavish et al., 2025c, a) is to also find rigorous explanations for DTs.

### v3

@@ Abstract
The Rashomon set of decision trees (DTs) finds several relevant uses. Recent work showed that DTs computing the same classification function, i.e. predictive equivalent DTs, can represent a significant fraction of the Rashomon set. Such redundancy can be undesirable. For example, feature importance based on the Rashomon set becomes inaccurate due the existence of predictive equivalent DTs, i.e. DTs that make the same prediction for every possible input. In recent work, McTavish et al. proposed solutions for several computational problems related with DTs, including that of deciding predictive equivalent DTs. The approach of McTavish et al. consists of applying the well-known method of Quine-McCluskey (QM) for obtaining minimum-size DNF (disjunctive normal form) representations of DTs, which are then used for comparing DTs for predictive equivalence. Furthermore, the minimum-size DNF representation was also applied to computing explanations for the predictions made by DTs, and to finding predictions in the presence of missing data. However, the problem of formula minimization is hard for the second level of the polynomial hierarchy, and there exist different reasons why the QM method may exhibit worst-case exponential running time and space. This paper first demonstrates that there exist decision trees that trigger the worst-case exponential running time and space of the QM method. Second, the paper shows that the QM method may incorrectly decide predictive equivalence, if two key constraints are not respected, and one may be difficult to formally guarantee. Third, the paper shows that any of the problems to which the minimum-size DNF representation has been applied to can in fact be solved in polynomial time, in the size of the DT. The experiments confirm that, for DTs for which the worst-case of the QM method is triggered, the algorithms proposed in this paper are orders of magnitude faster than the ones proposed by McTavish et al. Finally, the paper relates predictive equivalence with logic-based explanations and measures of importance.
@@ 1 Introduction
In decision tree learning, the Rashomon set (Breiman, 2001b; Xin et al., 2022) captures the set of decision trees (DTs) for a given classification problem which are within some distance from some accepted definition of optimal DT (Xin et al., 2022). Several uses of the Rashomon set have been studied in recent years (Fisher et al., 2019; Semenova and Rudin, 2019; Rudin et al., 2022; Semenova et al., 2022; Xin et al., 2022; Donnelly et al., 2023; Paes et al., 2023; Biecek et al., 2023; Kobylinska et al., 2024; Hsu et al., 2024; Ciaperoni et al., 2024; Li et al., 2024b, a; Donnelly et al., 2025; Nguyen et al., 2025). For the same classification problem, different decision trees capture different classification functions, and it is often of interest to study a sample of the decision trees in the Rashomon set. Recent work (McTavish et al., 2025c, a) studies predictive equivalent DTs in the Rashomon set, i.e. DTs which represent the same classification function. This earlier work observes that the predictive equivalent DTs in the Rashomon set can represent a fairly significant percentage of the DTs, and so their removal enables keeping in the Rashomon set only DTs that capture different classification functions, i.e. that are not predictive equivalent. For example, this yields better estimates in terms of feature importance (Donnelly et al., 2023).
@@
To identify DTs that are predictive equivalent, McTavish et al. (McTavish et al., 2025c, a) propose a solution that revisits work on boolean formula minimization from the 1950s, namely the well-known Quine-McCluskey (QM) boolean formula minimization method (Quine, 1952, 1955; McCluskey, 1956). [^1] There are different algorithms that implement variants of the QM method (McCluskey, 1956; Brayton et al., 1984; Coudert, 1994). The work of McTavish et al. (McTavish et al., 2025c, a) exploits the implementation of QM available in SymPy (Meurer et al., 2017). Starting from some DT $\mathcal{T}$, the QM method is applied to computing a minimum-size boolean formula $\mathcal{T}_{\mathsf{DNF},c}$ for each class $c$, in disjunctive normal form (DNF). This formula is then used not only for deciding predictive equivalence, but also to answer a number of important queries regarding DT prediction in terms of non-availability of some features and their explanation (McTavish et al., 2025c, a). It should be underlined that the formula ${\mathcal{T}}_{\mathsf{DNF},c}$ is not any DNF representation of a DT for predicting class $c$. Since the language of DTs is less expressive that that of DNF formulas (Hyafil and Rivest, 1976), then any DT can be represented by a DNF of size no larger that the DT. (Thus, this gives an upper bound on the smallest DNF size for representing a DT.) For each class $c$, ${\mathcal{T}}_{\mathsf{DNF},c}$ represents a minimum-size DNF formula obtained from the set of prime implicants for predicting that class. Moreover, ${\mathcal{T}}_{\mathsf{DNF}}$ represents the set of DNFs, one for each class $c$. The approach of McTavish et al. proposes the use of the QM method, subject to two constraints, to obtain a minimal representation, which is then used for answering different queries, including predictive equivalence.
@@
While it is known that the number of prime implicants can grow exponentially with the number of variables for arbitrary boolean formulas (Chandra and Markowsky, 1978), this paper shows that this is also the case with DTs, i.e. there exist DTs for which the BCF is worst-case exponential in the size of the DT. The paper then improves on earlier work (McTavish et al., 2025c, a), by proving that the computation of $\mathsf{BCF}_{c}(\mathcal{T})$ (which is worst-case exponential in size) and $\mathcal{T}_{\mathsf{DNF},c}$ (which requires solving an NP-hard optimization problem on a worst-case exponential size representation) are both unnecessary, and can be replaced with polynomial-time algorithms. Concretely, this paper proves that the properties of: i) completeness (Theorem 3.2 in (McTavish et al., 2025a)); ii) succinctness (Proposition 3.3 in (McTavish et al., 2025a)); and iii) the resolution of predictive equivalence (Theorem 3.4 in (McTavish et al., 2025a)) can all be solved in polynomial time in the size of the given DT. Furthermore, our paper underscores the formal correctness of all the proposed algorithms.
We also note that, in terms of practical implementations, the issues with McTavish et al.’s approach may extend beyond the inefficiency of the proposed algorithms (McTavish et al., 2025c, a). The use of the QM method for deciding predictive equivalence hinges on two constraints. First, the set of prime implicants (the Blake Canonical Form Blake (1937); Brown (1990)) must be sorted. Second, the execution of the set-covering phase of the QM method must be deterministic. These two conditions ensure that, even though a minimum-size DNF representation lacks canonicity, predictive equivalence can be correctly decided, and so are critical for correctness. Neverthless, even if the underlying algorithms are sequential and do not exploit non-determinism, it is the case that most programming languages do not ensure deterministic execution Mudduluru et al. (2021); Miao et al. (2025). Thus, and to the best of our knowledge, one of the conditions required my McTavish et al. is not formally guaranteed in practice.
@@
This paper is organized as follows. Section˜2 introduces the notation and definitions used throughtout. Section˜3 develops an in-depth analysis of the work of McTavish et al. (McTavish et al., 2025c, a). Section˜3.1 analyzes the approach of McTavish et al. approach in terms of running time, proving that the core algorithm can require exponential time and space, and also demonstrating that the worst-case running time and space is exercised in the case of DTs. Section˜3.2 goes a step further, and shows that the use of the QM method for deciding predictive equivalence can produce incorrect results without the constraints imposed by McTavish et al. Furthermore, it is argued that one of these constraints may be unrealistic to achive in practice. Section˜4 proposes polynomial-time algorithms for the three computational problems studied by McTavish et al. (McTavish et al., 2025c, a). Section˜6 proves a number of additional results, relating predictive equivalence with logic-based explanations (Marques-Silva, 2022; Marques-Silva and Ignatiev, 2022), but also with measures of importance, including the Shapley values (Shapley, 1953; Chalkiadakis et al., 2012) and Banzhaf values (Banzhaf III, 1965; Chalkiadakis et al., 2012). These additional results also allow developing an algorithm for predictive equivalence that asymptotically improves on the algorithm proposed in Section˜4. Section˜5 presents simple experiments demonstrating the massive difference in performance between the approach of McTavish et al. and the algorithms proposed in this paper. Section˜7 concludes the paper.
@@
Before this manuscript was submitted for publication, McTavish et al. (2025c, a) had been informed of the efficiency limitations of their proposed algorithms, and of the possible issues of soundness resulting from imposing the constraint of deterministic execution on the set-covering step of the QM method, since this constraint may be difficult to guarantee in practice Mudduluru et al. (2021).
@@ 2.1 Machine Learning (ML) Models
A classification problem is defined on a set of features ${\mathcal{F}}=\{1,\ldots,m\}$. Each feature $i$ has a domain $\mathbb{D}_{i}$. Domains can be categorical or ordinal. If ordinal, domains can be integer or real-valued. Feature space is defined as the cartesian product of the domains, $\mathbb{F}=\mathbb{D}_{1}\times\cdots\times\mathbb{D}_{m}$. A classifier computes a classification function $\kappa$, that maps feature space to a set of classes $\mathbb{K}=\{c_{1},\ldots,c_{K}\}$, i.e. $\kappa:\mathbb{F}\to\mathbb{K}$.
@@ 2.2 Logic Foundations
Let $\mathbb{OP}$ denote some universe of relational operators, e.g. we could consider $\mathbb{OP}=\mathbb{OP}_{s}\cup\mathbb{OP}_{t}$, with $\mathbb{OP}_{t}=\{\in\}$ and $\mathbb{OP}_{s}=\{=,<,\leq,\geq,>\}$. (Logic negation can be used to obtain additional operators.) Let $\mathbb{L}=\{(x_{i},R_{i},\mathsf{op}_{i})\,|\,(i\in{\mathcal{F}})\land(R_{i}\subseteq\mathbb{D}_{i})\land(\mathsf{op}_{i}\in\mathbb{OP}_{t})\}\cup\{(x_{i},R_{i},\mathsf{op}_{i})\,|\,(i\in{\mathcal{F}})\land(R_{i}\in\mathbb{D}_{i})\land(\mathsf{op}_{i}\in\mathbb{OP}_{s})\}$ denote all the literals that can be obtained from the variables associated with the features in $\mathcal{F}$} and subsets or values from their domains. (Throughout the paper, we will simplify the notation. For example, $(x_{i},v_{i})$ will be used instead of $(x_{i},v_{i},=)$, and both represent the literal $x_{i}=v_{i}$. Moreover, $(x_{i},v_{i},=)$ is equivalent to writing $(x_{i},\{v_{i}\},\in)$.) The set of partial assignments is defined by the powerset of $\mathbb{L}$, i.e. $\mathbb{A}=2^{\mathbb{L}}$. Thus, a partial assignment ${\mathcal{A}}\in\mathbb{A}$ is a subset of $\mathbb{L}$. In the case of complete assignments (or inputs to an ML model), we require that all features are assigned a single value from their domain, $\mathbb{I}=\{{\mathcal{A}}\in\mathbb{A}\,|\,\forall(i\in{\mathcal{F}}).|\{(x_{i},R_{i},\mathsf{op}_{i})\in{\mathcal{A}}\}|=1\land(x_{i},v_{i},=)\in{\mathcal{A}}\}$. With each literal $(x_{i},R_{i},\mathsf{op}_{i})\in\mathbb{L}$ we associate a domain $\mathsf{dom}((x_{i},R_{i},\mathsf{op}_{i}))=\{\mathbf{x}=(x_{1},\ldots,x_{i},\ldots,x_{m})\in\mathbb{F}\,|\,x_{i}\,\mathsf{op}_{i}\,{v_{i}}\}$. Thus, the domain of a literal is the set of points in feature space for which the literal holds true. Similarly, we define the domain of a partial assignment $\mathcal{A}$}, i.e. $\mathsf{dom}({\mathcal{A}})=\cap_{(x_{i},v_{i},\mathsf{op}_{i})}\mathsf{dom}(x_{i},v_{i},\mathsf{op}_{i})$. In addition, a partial assignment is said to be consistent, it if exhibits a non-empty domain:
$\mathsf{Consistent}({\mathcal{A}})\penalty 10000\ \coloneq[\mathsf{dom}({\mathcal{A}})\not$$\emptyset$]
By default, we will require partial assignments to be consistent. Finally, we note that (prime) implicants can also be represented as partial assignments (Dillig et al., 2012; Amgoud, 2023). This provides a convenient framework for formalizing explanations in DTs.
@@ 2.3 Classification with Decision Trees
A DT is a connected directed acyclic graph ${\mathcal{T}}=(V,E)$, with a single root node $r\in{V}$. $r=\mathsf{Root}({\mathcal{T}})$ indicates that $r\in{V}$ is the root of DT $\mathcal{T}$}. The tree nodes $V$ are partitioned into non-terminal and terminal nodes. Each non-terminal node $s$ (such that predicate $\mathsf{NonTerminal}(s;{\mathcal{T}})$ holds true) is associated with a feature $i$, i.e. $i=\mathsf{Feature}(s;{\mathcal{T}})$. For a non-terminal node $s$, associated with feature $i$, the domain of $i$ is partitioned among the children of node $s$. Each terminal node $s$ (such that predicate $\mathsf{Terminal}(s;{\mathcal{T}})$ holds true) is associated with a class from $\mathbb{K}$, denoted by $\mathsf{Class}(s;{\mathcal{T}})$. Whereas non-terminal nodes have two or more children, terminal nodes have no children. Moreover, for each node $s$ in the DT, $\mathsf{Parent}(s;{\mathcal{T}})$ represents the parent of $s$ in the tree. For the root node $r=\mathsf{Root}({\mathcal{T}})$, $\mathsf{Parent}(r;{\mathcal{T}})=\mathsf{none}$. As with any classifier, a DT computes a classification function $\kappa_{{\mathcal{T}}}$, mapping feature space to $\mathbb{K}$. For convenience, $\mathsf{NonTerminal}({\mathcal{T}})$ denotes the nodes in $V$ that are non-terminal. Similarly, $\mathsf{Terminal}({\mathcal{T}})$ denotes the nodes in $V$ that are terminal.
@@
Two DTs $\mathcal{T}$} and $\mathcal{T}$$^{\prime}$ are predictive equivalent if it holds that: $\forall(\mathbf{x}\in\mathbb{F}).\left(\kappa_{{\mathcal{T}}}(\mathbf{x})=\kappa_{{\mathcal{T}}^{\prime}}(\mathbf{x})\right)$. Thus, predictive equivalence signifies that the two DTs cannot be distinguished over any point in feature space. For any pair of DTs ${\mathcal{T}}_{1},{\mathcal{T}}_{2}$, the predicate $\mathsf{PredEquiv}({\mathcal{T}}_{1},{\mathcal{T}}_{2})$ takes value $\top$ (i.e. it holds true) if the pair of DTs is predictive equivalent; otherwise it takes value $\bot$. Although the paper focuses mostly on DTs, it is plain that the definition of predictive equivalence can be used to relate two arbitrary ML models. For example, we could have a DT that is predictive equivalent to some random forest (RF) (Breiman, 2001a). In Section˜6 we revisit predictive equivalence in this more general setting.
Although it is well-known that learning an optimal decision tree is NP-hard (Hyafil and Rivest, 1976), recent years witnessed a renewed interest in the learning of optimal decision trees (e.g. (Nijssen and Fromont, 2007; Bessiere et al., 2009; Nijssen and Fromont, 2010; Bertsimas and Dunn, 2017; Verwer and Zhang, 2017; Narodytska et al., 2018; Verwer and Zhang, 2019; Hu et al., 2019; Verhaeghe et al., 2020a; Avellaneda, 2020; Aglin et al., 2020a; Lin et al., 2020; Janota and Morgado, 2020; Hu et al., 2020; Aglin et al., 2020b; Verhaeghe et al., 2020b; Demirovic and Stuckey, 2021; Shati et al., 2021; Demirovic et al., 2022; Rudin et al., 2022; van der Linden et al., 2023), among many others), with the purpose of improving ML model interpretability. The rationale is that DTs are interpretable and so the model can be used for obtaining an explanation, e.g. the path consistent with the given instance (Molnar, 2020). Moreover, and although it is generally accepted that DTs are interpretable (Breiman, 2001b), [^4] The perception of DT interpretability can be traced back more than two decades. A quote from a 2001 paper by L. Breiman (Breiman, 2001b) illustrates this perception: “On interpretability, trees rate an A+”. there is no rigourous definition of interpretability or interpretable ML model and this is unlikely to change (Lipton, 2018). Furthermore, a number of recent works (Izza et al., 2020, 2022; Audemard et al., 2022b; Marques-Silva and Ignatiev, 2023; McTavish et al., 2025a) observe that even optimal decision trees need to be explained, e.g. using a formal definition of explanation (Marques-Silva, 2022; Marques-Silva and Ignatiev, 2022). Concretely, recent work demonstrated that if DTs are explained by reporting the path consistent with some instance, then the explanation can be arbitrarily redundant on the number of features (Izza et al., 2020, 2022). Accordingly, one of the concerns of the approach of McTavish et al. (McTavish et al., 2025c, a) is to also find rigorous explanations for DTs.